Use the discriminate to find the number of real solution to
x^2+3x-4=0

Answer by Alan3354(69443)

Use the discriminate to find the number of real solution to
x^2+3x-4=0
--------------
It's discriminant. Discriminate is a verb.
disc = b^2 - 4ac = 9 - 4*1*(-4) = 25
disc > 0 --> 2 real solutions